OpenStudy (anonymous):

Help me please! Evaluate the following expression using the values given: Find –a^2-3b^3+c^2+2b^3-c^2 if a =3, b=2 and c=-3. A)-67 B)-17 C)-31 D)1

OpenStudy (anonymous):

@Chachi123 Please help me.....

OpenStudy (dan815):

ah just replace a ,b,c with the numbers they give you

OpenStudy (dan815):

oops

OpenStudy (anonymous):

\[a^2-3 b^3+2 b^3+c^2-c^2=a^2-b^3 \]

OpenStudy (anonymous):

Okay thanks.......

OpenStudy (dan815):

careful dont forget that - infont of -a^2

OpenStudy (anonymous):

strange original presentation. @dan815 is correct. \[-a^2-3 b^3+2 b^3+c^2-c^2=-a^2-b^3 \]